# When Imitation Learning Outperforms Reinforcement Learning in Surgical Action Planning

### v1

@@ Abstract
Keywords:
Surgical Action Planning Imitation Learning Reinforcement Learning Temporal Planning Surgical AI
@@ 1 Introduction
The fundamental question for surgical AI systems is the optimal learning paradigm: should systems learn through imitation of expert demonstrations (IL) or through trial-and-error optimization via reinforcement learning (RL)? Teleoperated robotic surgery provides natural access to expert demonstrations, making IL attractive. However, RL could theoretically discover strategies beyond expert-level performance through exploration. Recent work in endovascular surgery has demonstrated RL’s potential for autonomous navigation tasks, with successful applications in mechanical thrombectomy achieving high success rates while maintaining safety constraints [4, 5]. More broadly, RL world models in particular have shown that single configurations with no hyperparameter tuning can outperform specialized methods across diverse benchmark tasks, complete farsighted tasks such as collecting diamonds in Minecraft without human data or curricula, and capture expectations of future events during autonomous driving [6, 7, 8].
@@ 2.3 Dual-task Autoregressive Imitation Learning (DARIL)
Our IL baseline models surgical action prediction as causal sequence generation, combining frame-level processing with autoregressive action generation:
@@
where $\mathcal{L}_{\text{current}}=-\sum_{t}\log p(a_{t}|f_{t-w+1:t})$ and $\mathcal{L}_{\text{next}}=-\sum_{t}\log p(a_{t+1}|f_{t-w+1:t})$ represent cross-entropy losses for direct prediction of the 100 action classes, $\mathcal{L}_{\text{embed}}=\sum_{t}||e_{t+1}-\hat{e}_{t+1}||^{2}$ is the MSE loss for next frame embedding prediction, and $\mathcal{L}_{\text{phase}}$ is the phase recognition loss.
@@ 4.2 Implications for Surgical AI
Bootstrapping RL with IL: Our findings suggest a promising approach for surgical AI: bootstrapping RL models with IL-learned basic skills, then using physics simulators or world models for safe exploration of new techniques. This addresses the fundamental challenge that exploration and mistakes are crucial for improving surgical techniques, but must be tested in simulation rather than on patients.
@@ 5 Conclusion
Future surgical AI development should carefully consider domain characteristics, data quality, and evaluation alignment when choosing between IL and RL approaches. While IL excels at expert behavior mimicking, RL’s exploration capabilities may prove valuable in comprehensive evaluation frameworks capturing patient outcomes beyond expert similarity.

### v3

# DARIL: When Imitation Learning outperforms Reinforcement Learning in Surgical Action Planning

@@ 1 Introduction
The fundamental question for surgical AI systems is the optimal learning paradigm: should systems learn through imitation of expert demonstrations (IL) or through trial-and-error optimization via reinforcement learning (RL) [2]? Teleoperated robotic surgery provides natural access to expert demonstrations, making IL attractive. However, RL could theoretically discover strategies beyond expert-level performance through exploration. Recent work in endovascular surgery has demonstrated RL’s potential for autonomous navigation tasks, with successful applications in mechanical thrombectomy achieving high success rates while maintaining safety constraints [16, 15]. More broadly, RL world models in particular have shown that single configurations with no hyperparameter tuning can outperform specialized methods across diverse benchmark tasks, complete farsighted tasks such as collecting diamonds in Minecraft without human data or curricula, and capture expectations of future events during autonomous driving [3, 4, 5].
@@ 2.3 Dual-task Autoregressive Imitation Learning (DARIL)
Our IL baseline follows an offline Behavior Cloning (BC) approach, using supervised learning on expert demonstrations to model surgical action prediction as causal sequence generation, combining frame-level processing with autoregressive action generation:
@@
[…]
@@ 4.2 Implications for Surgical AI
Bootstrapping RL with IL: Our findings suggest a promising approach for surgical AI: bootstrapping RL models with IL-learned basic skills, then using physics simulators or world models for safe exploration of new techniques, a strategy that aligns with emerging data-centric paradigms in the field [2]. This addresses the fundamental challenge that exploration and mistakes are crucial for improving surgical techniques, but must be tested in simulation rather than on patients.
@@ 5 Conclusion
Future surgical AI development should carefully consider domain characteristics, data quality, and evaluation alignment when choosing between IL and RL approaches. While IL excels at expert behavior cloning, RL’s exploration capabilities may prove valuable in comprehensive evaluation frameworks capturing patient outcomes beyond expert similarity.